# HLE-Verified: A Systematic Verification and Structured Revision of Humanity’s Last Exam

### v1

@@ 1 Introduction
As frontier language models (LMs) continue to advance rapidly, there is increasing demand for evaluation benchmarks that are simultaneously difficult, broad in disciplinary coverage, and resistant to saturation. High-difficulty benchmarks play a central role in tracking progress, informing capability claims, and shaping scientific discourse about model reasoning abilities. In this context, Humanity’s Last Exam (HLE) has emerged as a widely used benchmark for assessing model performance on challenging, multi-domain questions spanning mathematics, science, engineering, and the humanities Phan et al. (2025); for AI Safety Phan Long agibenchmark@ safe. ai 1 Gatti Alice 1 Li Nathaniel 1 Khoja Adam 1 Kim Ryan 1 Ren Richard 1 Hausenloy Jason 1 Zhang Oliver 1 Mazeika Mantas 1 Hendrycks Dan dan@ safe. ai 1 (2026). Results on HLE have been cited to support claims regarding reasoning depth, generalization capacity, and reliability of state-of-the-art models.
However, as benchmark difficulty increases, so does the importance of annotation integrity. When many items lie near a model’s decision boundary, even minor inconsistencies in problem statements or reference answers can disproportionately influence aggregate metrics. Recent independent audits and community-led examinations have raised concerns that HLE contains a non-trivial number of noisy items, including ambiguous or underspecified problem statements, incorrect reference answers, and inconsistencies between rationales and final answers A. White (2025); lhl (2026); 16. These analyses suggest that a portion of measured model performance on HLE may reflect annotation artifacts rather than genuine capability differences.
@@ 2.1 Why benchmark errors matter
Benchmarks serve as the primary measurement interface between model development and scientific claims about capability. Let a benchmark contain a fraction $\rho$ of flawed items. Even for small $\rho$, aggregate metrics such as accuracy or pass@k can be meaningfully distorted, particularly when flawed items are not randomly distributed. If such items are concentrated in specific domains, reasoning patterns, or difficulty strata, they may introduce systematic measurement bias Liang et al. (2022); Gema et al. (2025), favoring certain model families while penalizing others due to differences in training exposure, reasoning style, or calibration behavior.
Beyond aggregate accuracy, benchmark flaws pose a deeper challenge to interpretability. Many analyses assume a well-defined correspondence between correctness and model confidence, particularly in calibration- or uncertainty-aware evaluations Guo et al. (2017); Kadavath et al. (2022). When items are ill-posed, contain incorrect answer keys, or admit multiple valid interpretations, this correspondence becomes ill-defined: models may be penalized for producing valid but unanticipated answers, or rewarded for matching erroneous supervision. In such cases, apparent calibration error or ranking instability may reflect properties of benchmark noise rather than intrinsic model behavior Chao et al. (2024).
@@ 2.2 HLE as an evaluation substrate
However, prior work has demonstrated that large-scale benchmarks may contain systematic annotation errors or unstable evaluation artifacts that meaningfully affect model rankings Gema et al. (2025); Prathifkumar et al. (2025). To date, such concerns have not been systematically quantified for HLE, nor has their impact on evaluation outcomes been rigorously characterized. Consequently, it remains unclear to what extent observed performance differences on HLE reflect genuine capability gaps versus sensitivity to benchmark defects.

### v2

@@ 1 Introduction
As frontier language models (LMs) continue to advance rapidly, there is increasing demand for evaluation benchmarks that are simultaneously difficult, broad in disciplinary coverage, and resistant to saturation. High-difficulty benchmarks play a central role in tracking progress, informing capability claims, and shaping scientific discourse about model reasoning abilities. In this context, Humanity’s Last Exam (HLE) has emerged as a widely used benchmark for assessing model performance on challenging, multi-domain questions spanning mathematics, science, engineering, and the humanities (Center for AI Safety et al., 2026). Results on HLE have been cited to support claims regarding reasoning depth, generalization capacity, and reliability of state-of-the-art models.
However, as benchmark difficulty increases, so does the importance of annotation integrity. When many items lie near a model’s decision boundary, even minor inconsistencies in problem statements or reference answers can disproportionately influence aggregate metrics. Recent independent audits and community-led examinations have raised concerns that HLE contains a non-trivial number of noisy items, including ambiguous or underspecified problem statements, incorrect reference answers, and inconsistencies between rationales and final answers (A. White (2025); lhl (2026); 15). These analyses suggest that a portion of measured model performance on HLE may reflect annotation artifacts rather than genuine capability differences.
@@ 2.1 Why benchmark errors matter
Benchmarks serve as the primary measurement interface between model development and scientific claims about capability. Let a benchmark contain a fraction $\rho$ of flawed items. Even for small $\rho$, aggregate metrics such as accuracy or pass@k can be meaningfully distorted, particularly when flawed items are not randomly distributed. If such items are concentrated in specific domains, reasoning patterns, or difficulty strata, they may introduce systematic measurement bias (Liang et al., 2022; Gema et al., 2025), favoring certain model families while penalizing others due to differences in training exposure, reasoning style, or calibration behavior.
Beyond aggregate accuracy, benchmark flaws pose a deeper challenge to interpretability. Many analyses assume a well-defined correspondence between correctness and model confidence, particularly in calibration- or uncertainty-aware evaluations (Guo et al., 2017; Kadavath et al., 2022). When items are ill-posed, contain incorrect answer keys, or admit multiple valid interpretations, this correspondence becomes ill-defined: models may be penalized for producing valid but unanticipated answers, or rewarded for matching erroneous supervision. In such cases, apparent calibration error or ranking instability may reflect properties of benchmark noise rather than intrinsic model behavior (Chao et al., 2024).
@@ 2.2 HLE as an evaluation substrate
However, prior work has demonstrated that large-scale benchmarks may contain systematic annotation errors or unstable evaluation artifacts that meaningfully affect model rankings (Gema et al., 2025; Prathifkumar et al., 2025). To date, such concerns have not been systematically quantified for HLE, nor has their impact on evaluation outcomes been rigorously characterized. Consequently, it remains unclear to what extent observed performance differences on HLE reflect genuine capability gaps versus sensitivity to benchmark defects.